# The Surprising Effectiveness of Test-Time Training for Abstract Reasoning

### v1

@@ Abstract
[…]
@@ 1 Introduction
Large-scale neural language models (LMs) excel at performing tasks that occur in their training data, and often elementary variations or compositions of those tasks (Brown et al., 2020; Todd et al., 2024). Given natural language task specifications or a small number of examples, LMs often successfully infer the desired task and produce an appropriate output. But can LMs also solve new problems, involving non-trivial reasoning, planning, or string manipulation of a kind very different from their pre-training data? This question is central to understanding the novel skill acquisition capabilities of current AI systems, which has been proposed as a key measure of intelligence (Chollet, 2019).
@@
For complex and novel tasks, it is often difficult to obtain a correct answer simply by sampling from an LM (Wu et al., 2023). However, a significant finding in recent years has been that LM performance can be substantially improved by augmenting LM decoding with additional test-time computation. Methods in this category include chain-of-thought prompting (Wei et al., 2022), sampling with majority voting (self-consistency; Wang et al., 2022), code execution (Brown et al., 2024; Snell et al., 2024; Damani et al., 2024), and search (Yao et al., 2024).
One scaling strategy that has gained recent attention is test-time training (TTT), in which models are updated through explicit gradient steps based on test-time inputs (Krause et al., 2018; 2019). This method differs from standard fine-tuning as it operates in an extremely low-data regime—typically via an unsupervised objective on a single input, or a supervised objective applied to one or two in-context labeled examples. Modern versions of this approach was proposed for vision models by Sun et al. (2020), and also applied to sequence models by Gandelsman et al. (2022). The design space for TTT approaches is large, and there is currently a limited understanding of which design choices are most effective for LMs (and specifically for novel-task learning). In this paper, we systematically study the impact of various TTT design choices, as well as its interaction with pre-training and sampling schemes.
We evaluate these methods in the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019), a collection of extremely challenging few-shot visual reasoning problems. ARC is an ideal benchmark for testing the limits of LM generalization as it presents novel tasks, in a novel format, requiring nontrivial search and inference capabilities. Current language models perform poorly on ARC. Most successful approaches have relied on program synthesis techniques (Butt et al., 2024; Ainooson et al., 2023; Huang et al., 2023), though recently Cole et al. (2024) reported promising results using TTT on the benchmark.
We identify several crucial ingredients for effective application of TTT to few-shot learning: (1) initial fine-tuning on synthetic tasks similar to those encountered at test time, (2) an augmented, leave-one-out task generation strategy for constructing the test-time dataset, (3) per-instance adapter training and (4) a self-consistency (Wang et al., 2022) approach under invertible transformations. With careful choices of these components, TTT can significantly improve LM performance on ARC—increasing accuracy by up to a factor of six over a 1B model, and achieving state-of-the-art results for published, purely neural models on the ARC task with a 8B model. Indeed, our results show that when equipped with test-time training, ordinary LMs can match or exceed the performance of many neuro-symbolic approaches on ARC.
Our main contributions [^1] Our implementation can be found at this link. are:
@@
    We identify and systematically analyze the key components needed for test-time training on ARC tasks, with a a novel test time training data generation and self-consistency component.
@@
    We achieve state-of-the-art results among published neural approaches on the ARC validation set:
    • 53% accuracy on the public validation set with a 8B parameter model.
    • 61.87561.875% accuracy when ensembled with program synthesis approaches, comparable to average human performance on the dataset.
  3. 3.
    We demonstrate that tasks that could only be solved by program synthesis previously can be solved with fully neural approaches equipped with our TTT framework.
These results challenge the assumption that symbolic components are strictly necessary for solving such complex tasks. Instead, they suggest that the critical factor in solving novel reasoning problems may be the allocation of proper computational resources during test time, perhaps independently of whether these resources are deployed through symbolic or neural mechanisms.
@@ 2 Preliminaries
In this section, we first formally describe the ARC challenge. Next, we give an overview of in-context learning and test-time training, which form the foundation of our investigation. Finally, we detail our default experimental setup.
2.1 ARC Challenge
The Abstraction and Reasoning Corpus (ARC) aims to evaluate the abstract reasoning capabilities of language models through their ability to solve visual puzzles. Each puzzle, henceforth referred to as task, is comprised of input-output pairs of 2-D grids (up to $30\times 30$ in size) that contain shapes or patterns made with up to 10 different colors, as displayed in Fig.˜1(b). The output of each pair is obtained by applying an intuitive and shared transformation rule or function $y=f(x)$. In practice, these transformations are highly diverse and composite, ranging from simple concepts such as reflection and counting, to more complex ones such as application of gravity and path finding.
Each task in ARC is composed of a training and test split, with:
  • Training examples denoted $(x^{\textrm{train}}_{k},y^{\textrm{train}}_{k})_{k=1}^{K}$ (typically $K$ ranges from 2 to 7).
  • Test examples denoted $(x_{m}^{\textrm{test}},y_{m}^{\textrm{test}})_{m=1}^{M}$ (typically $M$ ranges from 1 to 3).
Given the set of training examples, the goal is to predict the test output $y^{\textrm{test}}$ for test test input $x^{\textrm{test}}$ by reasoning about the underlying transformation.
We denote a task as $d=(\mathbf{x}^{\textrm{train}},\mathbf{y}^{\textrm{train}},\mathbf{x}^{\textrm{test}},\mathbf{y}^{\textrm{test}})$ where $d\in\mathcal{D}_{\textrm{ARC}}$, the collection of such ARC tasks. The original training and validation sets of ARC dataset, respectively $\mathcal{D}^{\textrm{train}}_{\textrm{ARC}}$ and $\mathcal{D}^{\textrm{val}}_{\textrm{ARC}}$, consists of 400 tasks each. Success criteria requires to produce exact match for all test outputs (if not partial points are given). Please refer to Johnson et al. (2021) for a taxonomy and analysis of these tasks.
Most approaches to ARC can be categorized into two main categories: program synthesis and fully neural. Program synthesis approaches (Butt et al., 2024; Wang et al., 2024; Li et al., 2024; Greenblatt, 2024) try to first find the transformation function $f$, and later apply it to the test example. On the other hand, fully neural approaches (Thoms et al., 2023; Bober-Irizar and Banerjee, 2024) try to directly predict the output $y^{\textrm{test}}$, only implicitly reasoning about the underlying transformation. In this work, we use a fully neural approach, using a LM to predict the test outputs.
We start with an LM pre-trained on text data (without a vision encoder). To provide ARC examples as input to these models, we thus require a formatting function (denoted str) that converts 2D grids into their textual representations as shown in Fig.˜8. Previous work has presented examples as lists of numbers (Wang et al., 2024) or color words, or lists of connected components labeled with shapes and locations (Greenblatt, 2024). Given any such string representation of a task, we may present it to an LM and perform predictions with few-short prompting, as explained in the next section.
@@ 2.2 In-context Learning
At a certain scale, many LMs exhibit the ability to adapt to new tasks without updating their parameters by simply conditioning on input examples or instructions provided. Given a sequence of input-output pairs $(x_{1},y_{1}),...,(x_{n},y_{n})$ and a new input $x_{n+1}$, a LM can be used to generate the output $\hat{y}_{n+1}$ by sampling from:
$\hat{y}_{n+1}\sim\textrm{LM}(\cdot\mid x_{1},y_{1},\dots x_{n},y_{n},x_{n+1})$ (1)
The possibility of in-context learning as implicit machine learning simulation discussed in previous work (Akyürek et al., 2022), but the empirical evidence shows that in-context learning with language models does not always resemble any standard machine learning algorithm (Zhao et al., 2024; Min et al., 2022), and it does not always work out-of-the box for novel tasks — e.g. small language models (few billion parameters) performs poorly on ARC (Opielka et al., 2024; Bober-Irizar and Banerjee, 2024).
@@ 2.3 Test-Time Training
Figure 2: TTT dataset generation for a test task (Section˜3.1): We start by creating leave-one-out tasks from the given training examples of the task. These tasks are then augmented through rule-based transformations to obtain the full TTT dataset. Finally, we train task-specific LoRA adapters on top of the base FT model.
@@
Test-time training (TTT) enables parametric models to adapt during inference through dynamic parameter updates, an approach that remains relatively unexplored in the era of large language models. This technique is a form of transductive learning, where models leverages the test data structure to improve its predictions. The general TTT process works as follows: Starting with initial model parameters ${\bm{\theta}}_{0}$, for each test input (or batch of inputs), we first generate training data $\mathcal{D}_{\textrm{TTT}}(d_{\textrm{input}})$ from the test inputs. We then optimize these parameters to minimize a loss function $\mathcal{L}(\mathcal{D}_{\textrm{TTT}};{\bm{\theta}})$, producing temporarily updated parameters ${\bm{\theta}}_{d}$ for prediction. After generating predictions, the model is restored to the original parameters ${\bm{\theta}}_{0}$ for the next instance or batch. Thus, TTT trains a specialized prediction model for each test input, obtained by fine-tuning a base model on a test-time dataset generated from that test input.
In past work (e.g. Sun et al., 2020), $\mathcal{D}_{\text{TTT}}$ is typically constructed by applying an unsupervised objective (e.g. masked autoencoding) to the input $\mathbf{x}$ alone. However, the in-context learning setting we consider provides richer context in the form of demonstration pairs $(x_{1},y_{1}),\ldots,(x_{K},y_{K})$. Here, applying test-time tuning involves first constructing an initial language model $\mathrm{LM}$, mapping each test input $x$ to an input-specific dataset $\mathcal{D}_{\mathrm{TTT}}$, fine-tuning the LM to optimize some loss function $\mathcal{L}$ over the dataset according to: $\sum_{d\in\mathcal{D}_{\mathrm{TTT}}}\mathcal{L}(\mathrm{LM}(d))$, and finally sampling from the updated model to obtain a final prediction. Our experiments in this paper characterize each component of this pipeline, describing:
  1. 1.
    How to construct the augmented TTT dataset $\mathcal{D}_{\text{TTT}}$ from the test input (Section˜3).
  2. 2.
    An augmented inference strategy based on self-consistency over transformations (Section˜4).
  3. 3.
    A base model with parameters ${\bm{\theta}}_{0}$ that is fine-tuned on a dataset $\mathcal{D}_{\text{FT}}$ of similar tasks (Section˜5).
2.4 Experimental Setup
To investigate the impact of each TTT component, we conduct experiments by varying one component while holding the others constant at their optimal values (described in their respective sections). Our default configuration in the experiments uses the following settings:

### v2

# The Surprising Effectiveness of Test-Time Training for Few-Shot Learning

@@ Abstract
Language models (LMs) have shown impressive performance on tasks within their training distribution, but often struggle with structurally novel tasks even when given a small number of in-context task examples. We investigate the effectiveness of test-time training (TTT)—temporarily updating model parameters during inference using a loss derived from in-context examples—as a mechanism for improving LMs’ reasoning and few-shot learning capabilities. On the Abstraction and Reasoning Corpus (ARC), performing TTT with in-context examples yields up to 6×6$\times$6 × higher accuracy compared to fine-tuned baselines—reaching $53.0\%$ on the public validation set with an 8B-parameter LM and $61.9\%$ when ensembled with program-synthesis methods, matching average human performance. On BIG-Bench Hard (BBH), TTT on in-context examples surpasses standard few-shot prompting in the $10$-shot setting by $7.3$ percentage points ($50.5\%$ to $57.8\%$). Our findings highlight the limitations of in-context learning for novel tasks and demonstrate the potential of test-time training to enhance language model adaptability.
Machine Learning, ICML, Test-Time Training, In-Context Learning, Few-Shot Learning, ARC, BIG-Bench Hard
@@ 1 Introduction
Figure 1: Pass@2 accuracy on a subset of 80 randomly selected ARC validation tasks and overall accuracy on BIG-Bench Hard. The zero-shot baseline is 00 for ARC and $40.9\%$ for BBH, indicated by the dashed line. TTT boosts the performance of fine-tuned models (FT) on ARC by $27.5$ percentage points and increases accuracy on BBH by $7.3$ percentage points.
@@
Large-scale neural language models (LMs) have demonstrated remarkable success on few-shot learning of tasks related to those seen during pre-training, as well as elementary variations or compositions of those tasks (Brown et al., 2020; Todd et al., 2024). When given natural language specifications or a small number of examples, LMs can often infer the desired task and generate appropriate outputs. However, an open question is whether these models can truly acquire new skills for which they have not been trained—particularly, tasks involving non-trivial reasoning, planning, and abstraction in domains that differ significantly from their pre-training distributions. This question is fundamental to understanding how, and whether, LMs can exhibit the sort of flexible, novel-skill acquisition that has been proposed as a measure of intelligence (Chollet, 2019).
@@
Solving complex and novel tasks remains extremely challenging for LMs, and simple sampling approaches often yield poor performance on such problems (Wu et al., 2024; McCoy et al., 2024). However, recent progress has shown that LMs can be substantially improved by adding extra test-time computation. Several methods fall into this category, such as chain-of-thought prompting (Wei et al., 2022), sampling with majority voting (self-consistency; Wang et al., 2023), code execution (Brown et al., 2024; Snell et al., 2025; Damani et al., 2025), and search (Yao et al., 2023).
The idea of updating model parameters using instance-specific data at test time has roots in the literature on local learning (Bottou & Vapnik, 1992) and transductive learning (Joachims, 1999). In these methods, a learner refines its parameters or hypotheses after observing test inputs, adapting to individual examples or small clusters of examples. Such approaches inherently blur the line between training and inference, and can lead to robust adaptation in low-data scenarios or under distribution shift.
Modern versions of these transductive ideas for deep neural networks have been widely referred to as test-time training. In TTT, a model is updated at inference time using only the current test instance or a small batch of test instances, typically through explicit gradient steps. While test-time adaptation has been explored for vision models (Sun et al., 2020) and sequence architectures (Gandelsman et al., 2022; Sun et al., 2024; Behrouz et al., 2025), its interaction with other techniques for few-shot learning—especially in-context learning—remains less understood.
In this paper, we investigate how to leverage TTT on top of standard in-context learning (ICL) to boost performance on challenging tasks that require reasoning or rule-based generalization. In-context learning is a powerful means of adaptation without parameter updates, guided by short, task-specific prompts. We show that combining ICL with explicit gradient-based updates on test data can significantly improve performance on particularly difficult tasks. Specifically, our main contributions [^1] Code and data are available at https://github.com/ekinakyurek/marc (ARC) and https://github.com/adamzweiger/Fewshot-TTT (BBH). are:
@@
    A systematic analysis of the key components for effective test-time training, including strategies for selecting training data at inference, training objectives, and how TTT interacts with an LM’s pre-trained parameters and in-context learning.
@@
    An application of TTT to two challenging benchmark suites—The Abstraction and Reasoning Corpus (ARC; Chollet, 2019) and BIG-Bench Hard (BBH; Srivastava et al., 2023; Suzgun et al., 2023).
On ARC, our TTT approach outperforms existing open-source neural methods, attaining $53.0$% accuracy with an 8B model and $61.9$% when ensembled with a program-synthesis approach (comparable to human performance). On BBH, TTT yields a $7.3$% absolute improvement over few-shot prompting, achieving $57.8$% accuracy. Gains are particularly large on tasks involving structural rules or distribution shifts (e.g., Dyck languages, Ruin names), where TTT yields 20–50 percentage points of improvement over standard in-context prompting.
Overall, our findings highlight that TTT drastically improves LM’s few-shot learning ability on out-of-distribution tasks.
@@ 2.1 In-context Learning
At a certain scale, many LMs exhibit the ability to adapt to new tasks without updating their parameters by simply conditioning on input examples or instructions provided. Given a sequence of input-output pairs $(x_{1},y_{1}),\ldots,(x_{k},y_{k})$ and a new input $x_{k+1}$, an LM can generate the corresponding output $\hat{y}_{k+1}$ by sampling from:
$\hat{y}_{k+1}\sim\textrm{LM}(\cdot\mid x_{1},y_{1},\dots,x_{k},y_{k},x_{k+1})$
While the possibility of in-context learning (ICL) as implicit machine learning simulation is discussed in previous work (Akyürek et al., 2023), empirical evidence shows that in-context learning with language models does not always resemble standard machine learning algorithms (Zhao et al., 2024; Min et al., 2022b). Furthermore, ICL often struggles with novel tasks “out-of-the-box.” For example, large language models exhibit poor performance on datasets like ARC (Opiełka et al., 2024; Bober-Irizar & Banerjee, 2024).
@@ 2.2 Test-Time Training
Test-time training (TTT) enables parametric models to adapt during inference through dynamic parameter updates in response to each test input. This approach remains relatively unexplored in the era of large language models. The general TTT process is as follows: starting with initial model parameters ${\bm{\theta}}_{0}$, for each test input (or batch of inputs) $d$, we generate a temporary training dataset $\mathcal{D}_{\mathrm{TTT}}$. We then optimize these parameters to minimize a loss function
$\operatorname*{arg\,min}_{{\bm{\theta}}}\sum_{d_{\textrm{TTT}}\in\mathcal{D}_{\textrm{TTT}}}\mathcal{L}(\mathrm{LM}(d_{\textrm{TTT}};{\bm{\theta}})),$
resulting in temporarily updated parameters ${\bm{\theta}}_{d}$, which are subsequently used for prediction. [^2] Note that this use of “test-time training” is related but distinct from the one used in recent line of work wherein an RNN’s hidden state is treated as parameters and the update equation is interpreted as optimizing a recall-based regression objective (Ravi & Larochelle, 2017; Sun et al., 2024; Behrouz et al., 2025; Wang et al., 2025).
In previous work (e.g., Sun et al., 2020), $\mathcal{D}_{\mathrm{TTT}}$ is typically constructed by applying an unsupervised objective (e.g., masked autoencoding) to the input $x$ alone. In this paper, we extend TTT to the few-shot learning setting, treating it as a form of transductive learning by leveraging few-shot demonstration examples to improve predictions. Although TTT can also be applied to chain of thought (CoT; Wei et al., 2022), we focus on direct transduction, where demonstrations consist of input-output pairs $(x,y)$ without intermediate reasoning steps or explicit function descriptions.
The few-shot learning setting we consider provides richer context in the form of demonstration pairs $(x_{1},y_{1}),\ldots,(x_{K},y_{K})$. One simple method for TTT is Direct I/O training, where we directly treat each input-output $(x_{k},y_{k})$ pair as training instances. Our key insight is that the few-shot examples can also be used to construct a more robust and expansive $\mathcal{D}_{\mathrm{TTT}}$ of synthetic in-context learning tasks, allowing for effective model adaptation during test time. Additionally, when task-specific knowledge is available, this structure can be leveraged to further expand the dataset, as demonstrated in our experiments on ARC (Section 4). We also explore the general case where no task-specific information is used, as tested on BBH (Section 5).
Our experiments in this paper characterize each component of the TTT pipeline, investigating different design choices across the following stages: (1) constructing an input-specific training dataset $\mathcal{D}_{\mathrm{TTT}}$ at test-time; (2) fine-tuning the LM by optimizing a loss function $\mathcal{L}$ over the dataset $\sum_{d\in\mathcal{D}_{\mathrm{TTT}}}\mathcal{L}(\mathrm{LM}(d;{\bm{\theta}}))$; and (3) sampling from the updated model with an augmented inference strategy based on self-consistency to obtain a final prediction.
Figure 2: TTT design decisions. Data generation: A test task consists of input-output pairs $\{(x_{i},y_{i})\}$. The Leave-One-Out strategy removes one example at a time to form in-context learning tasks, while augmentations further expand the dataset. An alternative Direct I/O approach trains directly on the examples. Loss: The model is trained with loss computed on the Test Output (only the test-time prediction), All Outputs (including demonstration outputs), or Inputs and Outputs (all tokens). Parametrization: The Task-Specific approach trains a separate adapter per task while the Shared approach trains a single adapter across multiple tasks.
3 TTT Design
This section discusses the key design choices and challenges of applying TTT to LLMs, including how to best leverage their in-context learning capabilities, how to structure data for effective processing, what optimization objective to use, and how to efficiently update model parameters. We detail these considerations in the construction of the TTT dataset and the optimization setup (Figure 2).
3.1 Data Generation
Given a task with $K$ training input-output pairs $\{(x_{k},y_{k})\}_{k=1}^{K}$, a test-time training dataset $\mathcal{D}_{\textrm{TTT}}$ can be created by either following an in-context learning setup or a direct input-output (direct I/O) setup (top row in Figure 2):
Leave-one-out tasks
We begin with leave-one-out in-context learning tasks. For each pair $(x_{j},y_{j})$, we exclude it from the set of demonstrations and treat it as a “test” example within the newly formed synthetic task:
$d_{j}^{\textrm{ICL}}\;=\;\Bigl{(}\{(x_{k},y_{k})\}_{k\neq j},\,x_{j},\,y_{j}\Bigr{)}.$
Here, $\{(x_{k},y_{k})\}_{k\neq j}$ serves as the “in-context demonstrations,” and $(x_{j},y_{j})$ is the “synthetic test example.” To increase the number of synthetic tasks, we additionally permute the order of the demonstrations in each $d_{j}^{\textrm{ICL}}$.
Direct input-output (I/O) tasks
Rather than constructing in-context tasks, we treat each $(x_{k},y_{k})$ pair independently as a single training instance:
$d_{j}^{\textrm{I/O}}=(x_{j},y_{j})\,.$
In this setup, the model is fine-tuned on these training pairs without in-context demonstrations. While this approach is more computationally efficient, our results (Sections 4 and 5) show that it underperforms methods that utilize in-context demonstrations.
Data augmentation For certain tasks with structured inputs (e.g., ARC), we can apply invertible transformations (e.g., flips, rotations, color permutations) to further augment the TTT dataset. Let $\mathcal{T}$ be a set of invertible transformations. For each $t{\in}\mathcal{T}$, we have $t^{-1}(t(x))=x$, so we can apply $t$ to each training and test instance in $d_{j}$ to yield a transformed task $t(d_{j})$. Since these transformations preserve the core relationships in the data (e.g., the input-output pattern is the same, just rotated), they effectively expand the training signal. If rule-based transformations are used, the final TTT dataset is: $\mathcal{D}_{\textrm{TTT}}=\bigcup_{t\in\mathcal{T}}\bigcup_{j}t(d_{j}).$
3.2 Loss Function
We optimize the standard LM loss on $\mathcal{D}_{\textrm{TTT}}$. For the in-context leave-one-out setup, we experiment with $3$ different ways to take the loss (middle row in Figure 2):
  • Test output (no demonstration loss) The standard formulation where the loss is taken over $y_{\text{test}}$:
    $\mathcal{L}_{\textrm{LM}}^{\textrm{label}}=\mathcal{L}_{\textrm{LM}}(y_{\textrm{test}}\mid x_{1},y_{1},\ldots,x_{K},y_{K},x_{\textrm{test}};{\bm{\theta}})$
  • All outputs [^3] For ARC, we start the indexing at $k=2$ because the underlying transformation of an ARC task cannot be inferred without observing at least 1 demonstration. In addition to the loss on the test output, the loss is also taken over the outputs of the in-context demonstrations, which encourages the model to correctly predict the demonstration outputs after seeing the previous demonstrations:
    $\mathcal{L}_{\textrm{LM}}^{\textrm{outputs}}=\mathcal{L}_{\textrm{LM}}^{\textrm{label}}+\sum_{k=1}^{K}\mathcal{L}_{\textrm{LM}}(y_{k}|x_{1},y_{1},...,x_{k};{\bm{\theta}})$
  • Loss on inputs and outputs The loss is taken over all tokens, encouraging the model to learn the structure of $x$ as well as $y$:
    $\mathcal{L}_{\textrm{LM}}^{\textrm{all}}=\mathcal{L}_{\textrm{LM}}^{\textrm{outputs}}+\sum_{k=1}^{K}\mathcal{L}_{\textrm{LM}}(x_{k}|x_{1},y_{1},\ldots,y_{k-1};{\bm{\theta}})$
    This method, which requires learners to generate task inputs as well as outputs, is analogous to existing unsupervised TTT objectives (Sun et al., 2020).
We find in Sections 5.3 and 4.3 that the first method (taking the loss over both demonstration and test outputs) works best.
3.3 Parametrization
Once we have the test-time training dataset $\mathcal{D}_{\textrm{TTT}}$ (constructed via either the in-context or direct I/O approach), we perform a small number of gradient steps on task-specific LoRA adapters (Hu et al., 2022). This approach allows computationally efficient adaptation while maintaining the model’s general capabilities. By default, we learn task-specific LoRA adapters for each ARC or BBH task at test-time. That is, we obtain $K$ different LoRA adapters, where $K$ is the number of test tasks. We also experiment with using a single shared LoRA adapter from the aggregated dataset of few-shot examples drawn from multiple tasks (bottom row in Figure 2)—a test-time version of meta-ICL (Min et al., 2022a). We find that the shared adapter degrades performance on ARC, whereas it improves performance on BBH. We discuss this in more detail in Section 5.3.
4 Abstraction and Reasoning Corpus
4.1 Background
The Abstraction and Reasoning Corpus (ARC) aims to evaluate the abstract reasoning capabilities of language models through their ability to solve visual puzzles. Each puzzle (henceforth referred to as a task) consists of input-output pairs of 2D grids (up to $30\times 30$ in size) containing shapes or patterns in up to $10$ different colors, as displayed in Figure 3. The output of each pair is obtained by applying an intuitive and shared transformation or rule $y=f(x)$. Each task has 2-7 demonstration examples and 1-3 test examples.
Figure 3: Example of ARC and BBH tasks that the model successfully solves only after applying TTT.
4.2 Experimental Details